Tickets for a play at the community theater cost $12 for an adult and $8 for a child. If 130 tickets were sold and the total receipts were $1160, how many of each type of ticket were sold?
Answer by stanbon(48535)

Tickets for a play at the community theater cost $12 for an adult and $8 for a child. If 130 tickets were sold and the total receipts were $1160, how many of each type of ticket were sold?
--------------
Let # of adult tickets be "x" ; value of these is 12x dollars.
--------
# of child tickets is "130-x" ; value of these is 8(130-x)=1040-8x dollars
-----------
EQUATION:
value + value = 1160
12x + 1040-8x = 1160
4x = 120
x = 30 (# of adult tickets)
130-x = 100 (# of child tickets)
====================
Cheers,
Stan H.